# Parsing Through Boundaries in Chinese Word Segmentation

### v1

@@ Abstract
Yige Chen1∗    Zelong Li2∗    Changbing Yang3∗    Cindy Zhang3thanks: Equally contributed authors.    Amandisa Cady3 Ai Ka Lee3    Zejiao Zeng3    Haihua Pan1    Jungyeul Park3 1The Chinese University of Hong Kong, Hong Kong 2University College London, United Kingdom 3The University of British Columbia, Canada https://github.com/XDDz123/Chinese_WB_Dependency_Visualization
@@ 4.2 Systems
Figure 1: Screenshot of the visualization system using brat with the example sentence Gāi shū bèi fàngzhì zài xuéxiào túshūguǎn de fēi xiǎoshuō lèi wénxué zuòpǐn qū. (‘The book was placed in the nonfiction literature section of the school library.’). The visualization interface facilitates interactive exploration of Chinese dependency structures under different segmentation schemes. Users input sentences, select segmentation models via checkboxes, and choose a display mode—either grouping results by sentence or by model. Upon submission, the system generates corresponding dependency visualizations, enabling comparative syntactic analysis across segmentation strategies.
4.3 Experimental results
@@
4.4 Discussion
Our experimental results highlight the significant impact that word segmentation has on dependency parsing performance in Chinese. Variations in segmentation schemes lead to differences in the number and structure of tokens, which in turn affect both the syntactic representations learned by the parser and the evaluation metrics used to assess parsing quality. While stanza demonstrates stable parsing performance across most segmentation schemes, small but consistent fluctuations in UAS and LAS suggest that certain segmentation strategies may better align with the syntactic assumptions of the UD annotation framework.
The word segmentation F1 scores further reflect the challenge of reproducing original segmentation schemes using predictive models. Although stanza performs competitively, discrepancies between segmentation tools, such as LTP, CTB, and PKU, indicate differences in underlying linguistic assumptions about wordhood. These discrepancies underscore the importance of segmentation reproducibility when evaluating downstream tasks like dependency parsing.
In general, our findings confirm that segmentation is not a neutral preprocessing step in Chinese NLP but a linguistically consequential decision that shapes syntactic analysis. Effective parsing of Chinese text depends not only on the parsing model itself but also on the consistency and linguistic validity of the segmentation scheme used in training and evaluation.
@@ 5 Conclusion
This study presents a comparative analysis of dependency structures produced by different word segmentation schemes, offering insights into how segmentation strategies influence syntactic representations in Chinese NLP. Additionally, we develop a visualization tool designed to facilitate linguistic analysis and the debugging of models.
@@
This study focuses on Chinese dependency parsing using the GSD treebank and a limited set of word segmentation tools. While we demonstrate the influence of different segmentation strategies on parsing performance, our findings are constrained by the scope of models and datasets used. In particular, we rely on stanza for both segmentation and parsing, which may introduce model-specific biases. Furthermore, the evaluation does not incorporate human-annotated segmentation or parsing corrections beyond the GSD treebank, which may limit generalizability. Future work could extend this analysis to other treebanks, explore additional languages with similar challenges, and investigate segmentation-aware parsing models.
Ethic Statements
This work uses publicly available datasets and tools, including the Chinese GSD treebank from Universal Dependencies and open-source NLP frameworks such as stanza. No private or sensitive data were used in this study. The analysis focuses on linguistic structure and does not involve human subjects or personally identifiable information. All experiments were conducted in accordance with academic standards for reproducibility and transparency.

### v2

@@ Abstract
Yige Chen1,∗    Zelong Li2,∗    Cindy Zhang3,∗    Changbing Yang3,thanks: Equal contribution. Corresponding authors: Haihua Pan and Jungyeul Park.    Zejiao Zeng3 Amandisa Cady3    Ai Ka Lee3    Eunkyul Leah Jo3    Haihua Pan1,†    Jungyeul Park3,† 1The Chinese University of Hong Kong, Hong Kong 2University College London, United Kingdom    3The University of British Columbia, Canada https://github.com/XDDz123/Chinese_WB_Dependency_Visualization
@@ 4 Implementation and Discussion
Figure 1: Screenshot of the visualization system using brat with the example sentence Gāi shū bèi fàngzhì zài xuéxiào túshūguǎn de fēi xiǎoshuō lèi wénxué zuòpǐn qū. (‘The book was placed in the nonfiction literature section of the school library.’).
Experimental results
@@
Discussion
Our experimental results highlight the significant impact that word segmentation has on dependency parsing performance in Chinese. Variations in segmentation schemes lead to differences in the number and structure of tokens, which in turn affect both the syntactic representations learned by the parser and the evaluation metrics used to assess parsing quality. While stanza demonstrates stable parsing performance across most segmentation schemes, small but consistent fluctuations in UAS and LAS suggest that certain segmentation strategies may better align with the syntactic assumptions of the UD annotation framework. The word segmentation F1 scores further reflect the challenge of reproducing original segmentation schemes using predictive models. Although stanza performs competitively, discrepancies between segmentation tools, such as LTP, CTB, and PKU, indicate differences in underlying linguistic assumptions about wordhood. These discrepancies underscore the importance of segmentation reproducibility when evaluating downstream tasks like dependency parsing. In general, our findings confirm that segmentation is not a neutral preprocessing step in Chinese NLP but a linguistically consequential decision that shapes syntactic analysis. Effective parsing of Chinese text depends not only on the parsing model itself but also on the consistency and linguistic validity of the segmentation scheme used in training and evaluation.
@@ 5 Conclusion
This study presents a comparative analysis of dependency structures produced by different word segmentation schemes, offering insights into how segmentation strategies influence syntactic representations in Chinese NLP. Broader implications of Chinese word segmentation can also be explored in future work. [^6] For example, see Appendix B. Additionally, we develop a visualization tool to support linguistic analysis and model debugging.
@@
This study focuses on Chinese dependency parsing using the GSD treebank and a limited set of word segmentation tools. Our assumption is that GSD reflects the finest segmentation granularity among the available resources, and we construct coarser granularities by merging its tokens accordingly. While this allows for a controlled comparison across segmentation strategies, the assumption may not hold for other treebanks in Universal Dependencies, which often adopt different segmentation guidelines. As such, our findings are constrained by the scope of models and datasets used. Future work could extend this analysis to other treebanks, explore additional languages with similar challenges, and investigate segmentation-aware parsing models.
Ethics and Broader Impact Statement
This work uses publicly available datasets and tools, including the Chinese GSD treebank from Universal Dependencies and open-source NLP frameworks such as stanza. No private or sensitive data were used. The analysis focuses solely on linguistic structure and does not involve human subjects or personally identifiable information. All experiments were conducted in accordance with academic standards for reproducibility and transparency.
@@ Appendix A Licenses
Appendix B GEC Annotation under Different Word Boundary Schemes
Grammatical error correction is a downstream task that is highly sensitive to word segmentation choices, particularly in Chinese, where word boundaries are not overtly marked. Our extended analysis illustrates how different segmentation schemes shape not only tokenization but also the annotation of learner language.
Word boundaries play a crucial role in Chinese language processing and education, yet their ambiguity presents challenges for both natural language processing systems and language learners. This paper investigates the impact of WB segmentation on grammatical error annotation in L2 Chinese writing, aiming to identify the most effective WB strategies for supporting learners and instructors. We examine various segmentation approaches, comparing outputs produced by different NLP models to assess their influence on error detection and classification.
We refine the Chinese errant framework by incorporating a new error typology that accounts for phonetic, visual, and structural similarities in character misuses. Additionally, we adapt word-based segmentation schemes to improve annotation accuracy, mitigating tokenization inconsistencies found in the Chinese GSD treebank. A perception experiment involving L2 learners and instructors further evaluates how WB variations affect grammatical error identification and pedagogical effectiveness.
Our primary implementation builds on the jp-errant framework for Chinese GEC (Wang et al., 2025), which performs edit-based alignment and error categorization using part-of-speech tags and token-level comparisons. By default, jp-errant employs stanza (Qi et al., 2020) with GSD-style segmentation, which adopts a fine-grained, morpheme-oriented notion of wordhood. This approach captures subtle syntactic and morphological errors but can also fragment multiword expressions, complicating certain types of error interpretation.
To enable a more direct comparison with prior work such as ChERRANT (Zhang et al., 2022), which is based on the LTP analyzer (Che et al., 2010), we provide a stanza-based reimplementation of LTP-style segmentation. In contrast to GSD, the LTP scheme prioritizes lexical cohesion, grouping compounds and fixed expressions into single tokens. This often aligns more closely with learner intuitions and yields clearer boundaries for identifying lexical and syntactic errors in educational settings.
These findings suggest that word segmentation is not merely a preprocessing step but a structural choice that significantly influences error interpretation and system evaluation. The broader implications of segmentation granularity extend to annotation protocols, evaluation metrics, and the design of learner-aware NLP tools for Chinese.
Having both segmentation variants allows us to isolate the role of word boundaries in shaping annotation outcomes. For example, under GSD-style segmentation, a compound like {CJK}UTF8gbsn中山南路 Zhōngshān Nánlù (‘Zhongshan South Road’) may generate multiple edit operations if only one character is erroneous. In contrast, under LTP-style segmentation, the entire compound may be treated as a single token, yielding a different classification. These structural differences affect both alignment and labeling, ultimately influencing how system performance is assessed. By supporting both GSD-style and LTP-style segmentation in our GEC annotation pipeline, we promote more robust and interpretable evaluations of GEC systems for Chinese.